OpenStudy (anonymous):

How to simplify the expression: ((1/((x+h)^1/2)-(1/(x^1/2)))/h

OpenStudy (anonymous):

it seems u r finding the derivative of 1/sqrt(x) using limits

OpenStudy (anonymous):

first take LCM and then multiply by something like sqrt(x) +sqrt(x+h) both numerator and denominator and then ytaking limits will be easier as h may cancel out

OpenStudy (anonymous):

The section I doing practice for doesn't involve derivatives... It's only requiring me to simplify it algebraically, any suggestions?

OpenStudy (anonymous):

no need to worry its same may be uare in the pre-calculus stage...

OpenStudy (anonymous):

Could you walk me through on how to do it?

OpenStudy (anonymous):

ur question simplifies to (sqrt(x) - sqrt(x+h) )/(h *sqrt(x)*sqrt(x+h)) =(sqrt(x) - sqrt(x+h) ) * (sqrt(x) - sqrt(x+h) )/ (h *sqrt(x)*sqrt(x+h) *(sqrt(x) - sqrt(x+h) ) ) =(x-x-h)/ (h *sqrt(x)*sqrt(x+h) *(sqrt(x) - sqrt(x+h) ) ) = -h/ (h *sqrt(x)*sqrt(x+h) *(sqrt(x) - sqrt(x+h) ) ) =-1 (sqrt(x)*sqrt(x+h) *(sqrt(x) - sqrt(x+h) ) )

OpenStudy (anonymous):

Thank you so much, that really helped with the breakdown!

OpenStudy (anonymous):

The program I'm using said that the answer wasn't recognized..

OpenStudy (anonymous):

there may be extra left or right braces ..